two cars start at the same place and time, and travel in opposite directions. One car is traveling 15kph faster than the other. After 5 hours, the two cars are 275 km apart. Find the speed of each car.
Answer by rfer(12730)

x+x+15=275/5
2x+15=55
2x=40
x=20 mph
x+15=35 mph